if a force of 6900N is raised in a height of 15meters in 36 seconds what is the work done and the power developed?


1
Expert's answer
2022-05-06T01:48:03-0400

The work done

"W=Fd=6900*15=103500\:\rm J"

The power developed

"P=\frac{W}{t}=\frac{103500}{36}=2875\:\rm W"